What percent of 72 is 18? Use proportions.
Found 2 solutions by stanbon, ad_alta:
Answer by stanbon(75887)

What percent of 72 is 18? Use proportions.
------------------
x/100 = 18/72
x = 1800/72 = 25
--------------------
Ans: 25%
====================
Cheers,
Stan H.
=============

Answer by ad_alta(240)

1) 18/72=x/100
2) x=100*(18/72)
43) x=25%